### v1

# Deformable Mamba for Wide Field of View Segmentation

Abstract
Wide-FoV cameras, like fisheye and panoramic setups, are essential for broader perception but introduce significant distortions in 180° and 360° images, complicating dense prediction tasks. For instance, existing MAMBA models lacking distortion-aware capacity cannot perform well in panoramic semantic segmentation. To address this problem, this work presents Deformable Mamba, a unified framework specifically designed to address imaging distortions within the context of panoramic and fisheye semantic segmentation. At the core is a decoder constructed with a series of Deformable Mamba Fusion (DMF) blocks, making the whole framework more deformable, efficient, and accurate, when handling extreme distortions. Extensive evaluations across five datasets demonstrate that our method consistently improves segmentation accuracy compared to the previous state-of-the-art methods tailored for specific FoVs. Notably, Deformable Mamba achieves a +2.5% performance improvement on the 360° Stanford2D3D dataset, and shows better results across FoVs from 60° to 360°.
Comparison of wide Field of View (wide FoV) segmentation.
Omni-directional performance across wide FoV from 60° to 360°.
Figure 1: Wide Field of View (wide-FoV) segmentation is investigated on 180° and 360° images. (a) Compared to other methods tailored for 90° or 180° data, our unified Deformable Mamba framework generalizes well on wide-FoV segmentation. (b) Besides, our methods yield consistently better results (mIoU %) across all different FoVs as compared to the baseline on the 360° Matterport3D dataset.
11footnotetext: Equal contribution.22footnotetext: Corresponding author (e-mail: jiaming.zhang@kit.edu).
1 Introduction
Dense image analysis like image segmentation [2, 58, 21] is a fundamental task in computer vision, forming a crucial component of numerous downstream vision-based applications [70]. At the same time, advancements in sensor technology have led to a variety of sensor types, each with unique characteristics that impact imaging. In recent years, diverse sensors such as pinhole cameras, fisheye cameras [64, 48] with 180° field of views (FoV), and even panoramic cameras [3, 1, 67] with 360° FoV have been increasingly adopted across vision tasks. However, differences in camera construction and sensor type result in distinct imaging characteristics, which can complicate traditional image processing [66]. Adapting algorithms to perform effectively with a specific FoV often requires considerable time and resources, involving method redesign, parameter tuning, and model retraining to address the unique challenges posed by each camera type.
Based on our experiments (Fig. Deformable Mamba for Wide Field of View Segmentation), we observe that straightforwardly applying or transferring existing specific models to new sensor types frequently leads to significant performance degradation. For instance, methods [21] specifically designed for narrow-FoV pinhole cameras typically cannot generalize well when used for analyzing fisheye images with 180° FoV. This is primarily due to the high degree of image distortion and object deformation present in wide-FoV cameras, which complicates model generalization. Standard methods struggle to accurately capture the diverse object shapes and spatial distributions caused by such deformations. For instance, as shown in Fig. 2, the altered pixel arrangement in wide-FoV images might lead to misclassifications and limits their adaptability. Previous methods have proposed solutions targeting specific sensor types, such as CNN-based and transformer-based semantic segmentation models [52, 31] optimized for fisheye images, as well as approaches [66, 67] tailored to mitigate distortion in panoramic segmentation. So far, to address diverse distortions across different camera types in a unified manner remains under-explored.
Figure 2: Comparison of camera imaging with pinhole (left) and fisheye (right) cameras with different wide-FoV. Pinhole cameras maintain geometric fidelity but limited coverage, whereas wide-FoV cameras offer expansive scene capture while introducing substantial geometric distortions.
To address these challenges, we propose the Deformable Mamba model (DMamba for short), a versatile solution for image segmentation across varying wide-FoVs. In this work, a unified framework is created to effectively handling images from multiple camera types and sources, spanning FoVs from 180° to 360°, from indoor to outdoor scenarios, in synthetic and real-world imageries. At the core of our approach is a novel deformable decoder, which synergistically integrates State Space Models (SSMs) with deformable convolutions. Specifically, we leverage quadri-directional selective scan operations derived from VMamba [37] to capture global contextual information, while simultaneously employing deformable convolutions [77] to enhance feature representation. This dual-pathway desin effectively addresses the limitations of scanning mechanisms of Mamba [16] in visual applications while also endowing the model with the ability to perceive distortion. Our deformable decoder accommodates varied receptive fields, enabling seamless integration with CNN-based, transformer-based, and Mamba-based architectures. By leveraging this unified decoder, Deformable Mamba effectively captures features across disparate imaging perspectives, making it an adaptable solution for segmentation tasks across a broad spectrum of camera types.
To evaluate the effectiveness of Deformable Mamba, we conducted extensive experiments across five datasets for semantic segmentation, including three 360° (SynPASS [67], Stanford2D3D [1], Matterport3D [3]) and two 180° (WoodScape [64], SynWoodScape [48]) datasets. The data encompass a variety of scenarios including 180° and 360° images, indoor and outdoor environments, and both synthetic and real-world data. Our unified framework performs consistently well across various datasets. Notably, on 360° Stanford2D3D dataset, Deformable Mamba achieved state-of-the-art performance, with a ${+}2.5\%$ absolute improvement over the pinhole Mamba model. In 180° SynWoodScape dataset, it similarly surpassed prior best-performing models, achieving an increase of ${+}2.2\%$ in mean Intersection over Union (mIoU). Further ablation studies, including comprehensive omni-directional comparisons, consistently demonstrated the model’s effectiveness in handling wide-FoVs, showing its potential as a general solution for dense image analysis across diverse sensor types. We hope this work opens up new possibilities for unifying wide-FoV segmentation.
To summarize, our contributions are as follows.
  • We propose a Mamba-based framework termed Deformable Mamba, with the aim at unifying distortion-aware wide-FoV segmentation.
  • A deformable decoder constructed with Deformable Mamba Fusion (DMF) module is specifically designed to boost the performance of mamba-based models in the wide-FoV segmentation task.
  • Extensive experiments on five datasets across a variety of indoor and outdoor, synthetic and real-world wide-FoV scenes, demonstrate the state-of-the-art performance of our method in wide-FoV segmentation.
2 Related Work
2.1 Wide Field of View Segmentation
Wide Field-of-View (Wide-FoV) semantic segmentation has received increasing attention as it enables broader scene understanding in applications such as autonomous driving and robotics. However, directly applying pinhole models [39, 2, 5, 15, 26, 65, 65, 36, 34, 72, 50, 58, 9, 19, 6, 70] to wide-FoV images often results in downgraded performance due to object deformations and image distortions. Previous work [61, 62] propose the Panoramic Annular Semantic Segmentation (PASS) framework using a panoramic annular lens system.  [27] introduces panoramic panoptic segmentation for panoramic views. In [66, 67], deformable modules are proposed to enhance the distortion perception capability in transformers. Zheng et al. [71] introduce the Open Panoramic Segmentation (OPS) task in an open-vocabulary context. Apart from panoramas, previous methods [31, 11] revolve around multi-task on fisheye images, while a pretraining paradigm [42] and semi-supervised approaches [45, 43] are introduced to address the challenges of fisheye data. Manzoor et al. [40] substitutes the conventional convolutional modules in U-Net [47] and residual U-Net [46] with restricted deformable convolution (RDC) [11], imparting fisheye-distortion awareness to the model. Departing from these approaches, we are the first to propose a unified model applicable to all wide-FoV inputs, including panoramic and fisheye images.
2.2 State Space Model
CNNs [30] have achieved remarkable success in computer vision through effective local feature extraction and hierarchical representation learning capabilities, while Transformers [54] excel at modeling long-range dependencies despite their quadratic computational complexity. Recently, the state-space-model (SSM) based Mamba [16] architecture has emerged as a promising alternative, offering linear complexity and parallel computation through its selective scan mechanism, ushering in a new paradigm for long sequence modeling. The emergence of Mamba in sequence modeling has naturally extended to vision tasks, Zhu et al. [75] adopt the patch embedding from Vision Transformer (ViT) [12] to project 2D image patches into sequential tokens compatible with Mamba architecture. To address the inherent discrepancy between image and sequence representations, a quadri-directional scanning and merging mechanism termed SS2D is introduced in [37] to supersede Mamba’s canonical 1D scan operation. Subsequent work [25, 44, 24, 32, 69, 63, 4] introduce various scanning methods to explore variations of Mamba in visual representations. However, the study of using SSM for wide-FoV imagery is under-explored. Therefore, for the first time, we propose deformable Mamba to address both panoramic and fisheye semantic segmentation.
3 Methodology
3.1 Preliminary
Image Distortion. As shown in the left side of Fig. 2, the point $A$ in the world coordinates undergoes extrinsic transformation to camera coordinates by Eq. (1), where $R$ and $T$ denote the rotation matrix and translation vector, respectively. Following Eq. (2), points in the camera coordinates are projected onto the image plane, with $f_{x}$ and $f_{y}$ representing the focal lengths and $c_{x},c_{y}$ defining the principal point. While straight lines are preserved in pinhole-approximate narrow-FoV cameras, only points within the sensor’s FoV (red box in Fig. 2) are imaged, excluding points like $B^{\prime}$ that fall outside this region. Narrow-FoV cameras follow perspective projection principles where object size in the image varies inversely with distance between object and camera, introducing moderate distortion.
$\begin{bmatrix}\mathbf{X}_{c}\\ \mathbf{Y}_{c}\\ \mathbf{Z}_{c}\end{bmatrix}=\mathbf{R}\begin{bmatrix}\mathbf{X}_{w}\\ \mathbf{Y}_{w}\\ \mathbf{Z}_{w}\end{bmatrix}+\mathbf{T},$ (1)
$\begin{bmatrix}\mathbf{x^{\prime}}\\ \mathbf{y^{\prime}}\\ 1\end{bmatrix}=\begin{bmatrix}\mathbf{f}_{x}&0&\mathbf{c}_{x}\\ 0&\mathbf{f}_{y}&\mathbf{c}_{y}\\ 0&0&1\end{bmatrix}\begin{bmatrix}\frac{\mathbf{X}_{c}}{\mathbf{Z}_{c}}\\ \frac{\mathbf{Y}_{c}}{\mathbf{Z}_{c}}\\ 1\end{bmatrix}.$ (2)
In contrast, wide-FoV cameras like fisheye cameras employ complex projection models involving light refraction. Kannala et al. [28] propose a generalized projection model for wide-FoV cameras, the distance between the image point and the principal point $r$ is approximated by:
$\mathbf{r}(\mathbf{\theta})=\mathbf{k_{1}}\mathbf{\theta}+\mathbf{k_{2}}\mathbf{\theta}^{3}+\mathbf{k_{3}}\mathbf{\theta}^{5}+\mathbf{k_{4}}\mathbf{\theta}^{7}+\mathbf{k_{5}}\mathbf{\theta}^{9},$ (3)
where $\theta$ is the angle between the principal axis and the incoming ray, $k_{i},i\in\{1,2,3,4,5\}$ are the parameters of wide-FoV camera model. While these cameras capture wider FoV, they suffer from more severe geometric distortions compared to the perspective projection of narrow-FoV cameras. As shown in the right side of Fig. 2, barrel distortion in wide-FoV cameras causes straight lines to appear curved outward, particularly pronounced near image edges while the center remains relatively undistorted. This distortion manifests as radially-increasing displacement from the center, resulting in progressive spatial compression toward the image periphery.
State Space Model. State space model (SSM), which describes the dynamics of a system, can be formulated as:
$\begin{split}\mathbf{h}^{\prime}(\mathbf{t})&=\mathbf{A}\mathbf{h}(\mathbf{t})+\mathbf{B}\mathbf{x}(\mathbf{t}),\\ \mathbf{y}(\mathbf{t})&=\mathbf{C}\mathbf{h}(\mathbf{t}),\\ \end{split}$ (4)
where $x(t)$ and $y(t)$ denote the inputs and outputs of the system, respectively, $h(t)$ is the inner states of the system, $A,B,C$ are the system parameters. Gu et al.  [18] propose Structured State Space Sequence Model (S4) by initializing the parameters $A$ using Hippo Matrix (5) and discretizing parameters A and B by step size $\Delta$.
$\mathbf{A}_{nk}=-\begin{cases}(2\mathbf{n}+1)^{1/2}(2\mathbf{k}+1)^{1/2}&\text{if }\mathbf{n}>\mathbf{k}\\ \mathbf{n}+1,&\text{if }\mathbf{n}=\mathbf{k}\\ 0,&\text{if }\mathbf{n}<\mathbf{k}.\end{cases}$ (5)
In Mamba [16], a selective scan mechanism is proposed to further extend the SSM from time-invariant to time-variant, while also enabling parallel training. Observations. Based on HiPPO theory [17], Mamba [16] is inherently a sequence model, where the output of the current token only depends on preceding tokens, without considering subsequent ones - a limitation that becomes problematic for 2D images with inherent spatial structures. Unlike Transformer’s [54] mechanism in ViT [12], where attention computation for each patch considers all patches globally, Mamba processes image patches sequentially.
To address this limitation, previous work [37, 75, 4, 76, 63] have explored different scanning patterns, including bidirectional or quadri-directional scanning approaches, partially addressing the challenges posed by Mamba’s unidirectional scanning nature in visual tasks. However, these multi-directional scanning approaches face another challenge. As illustrated in Fig 3., consider an input image embedded into 9 patches (labeled 1-9). When modeling patch 5 using SS2D’s [37] quadri-directional scanning mechanism, the scanning sequences are:
  • (1) Left-Right: 1 - 2 - 3 - 4 - 5
  • (2) Right-Left: 9 - 8 - 7 - 6 - 5
  • (3) Top-Down: 1 - 4 - 7 - 2 -5
  • (4) Down-Top: 9 - 6 - 3 - 8 - 5
Figure 3: Visualization of SS2D quadri-directional scanning.
Patch 1 and patch 3 maintain equal distances from patch 5, demonstrating uniform spatial distribution. As illustrated in sequences (1) (3) (4), the scanning trajectory from patch 1 to patch 5 traverses double the path length compared to the direct route between patch 3 and patch 5, despite their equivalent spatial distances. Since the Mamba model learns relationships between serialized tokens, the contribution of earlier patches in the scanning path to the current patch decreases as their distance from the current patch increases. Thus, along this scanning path, we infer that patch 1 is more important to patch 5 than patch 3. This priority is introduced by the scanning mechanism, but it is not a reasonable assumption. Additionally, when computing for other target patches, this priority changes based on the target’s position.
Figure 4: Overview of the Deformable Mamba (DMamba) framework. Given wide-FoV images (180° or 360°), the features extracted by an encoder, are fused by the proposed Deformable Mamba Decoder, which constructed by four Deformable Mamba Fusion (DMF) modules.
3.2 Overall Architecture
An overview of our Deformable Mamba architecture is presented in Fig. 4, which illustrates how we make Mamba-based models deformable and available for wide-FoV semantic segmentation. It mainly follows the Encoder-Decoder architecture. The backbone selection is flexible for different methods like CNN-, transformer- or Mamba-based architectures. In this work, we explore with VMamba [37] as the backbone and our proposed Deformable Mamba Decoder. In our framework, there are two model sizes: DMamba-M (Mini) and DMamba-T (Tiny).
More specifically, given the wide-FoV input image in the shape of $H\times W\times 3$, a stem module is firstly utilized to patchify the original image into the shape of $\frac{H}{4}\times\frac{W}{4}\times 3$, the encoder gradually down-samples feature maps in the $l^{th}$ stage with channel dimensions $C_{l}\in\{96,192,384,768\}$ and resolutions $\in\{\frac{H}{4}\times\frac{W}{4},\frac{H}{8}\times\frac{W}{8},\frac{H}{16}\times\frac{W}{16},\frac{H}{32}\times\frac{H}{32}\}$. The features extracted by the encoder undergo a hierarchical fusion process in the decoder.
3.3 Deformable Mamba Decoder
The Deformable Mamba Decoder comprises four fusion and upsampling modules, where we introduce a novel Deformable Mamba Fusion (DMF) module for multi-scale feature integration. The initial DMF block processes dual inputs of ${\frac{H}{32}\times\frac{H}{32}\times 768}$ feature maps derived from the final encoder layer. The fused features are then processed through an upsampling module that simultaneously reduces channel dimensionality and increases spatial resolution, followed by channel mixing via an MLP layer. Following [49, 68, 35, 60], we utilize Pixel Shuffle [49] as the upsampling operation. In the subsequent stages, the second through fourth DMF blocks implement skip connections, facilitating interaction between the previously fused and upsampled features and their corresponding encoder features at matching scales. This progressive fusion strategy enables comprehensive multi-scale feature integration. The decoder’s output undergoes final processing through a fully connected layer and class mapping to generate the segmentation predictions.
3.4 Deformable Mamba Fusion
To address the aforementioned limitations of Mamba’s scanning mechanism in visual tasks (Sec. 3.1), we propose the DMF module, which simultaneously adapts the features extracted by the Mamba-based encoder and integrates multi-scale features. As illustrated in the lower part of Fig. 4, the DMF module takes two inputs: one is the feature $\mathbf{D}_{j}$ with dimensions $c\times h\times w$ output from the upper decoder layer, and the other is the corresponding encoder features $\mathbf{E}_{i}$ at matching scales, where $i$ and $j$ satisfy the relationship $i+j=5$. We employ the Cross-Scan [37] that scans $\mathbf{D}_{j}$ in four directions: (1) Left to Right, (2) Right to Left, (3) Top to Down, and (4) Down to Top, utilizing Mamba model’s global perception to generate four feature maps with resolution $c\times h\times w$ Following Cross-Merge [37], we first reverse features (2) and (4) along their last dimension and add them to (1) and (3) respectively, then transpose the $h$ and $w$ dimensions of feature (3) before adding it to (1).
Inspired by previous works [10, 77, 56, 59], we propose a simple yet effective approach, utilizing deformable convolution [77] to enhance the features extracted by the Mamba model. Specifically, for a feature point in the encoder feature map $\mathbf{E}_{i}$, we attend to $\mathbf{K}$ neighboring feature points, allowing the model to learn offsets that compensate for the biases introduced by the scanning process. Additionally, convolution operations on these selected regions provide a more localized attention mechanism, further enhancing the visual representation capability of Mamba-based models in visual tasks. The mathematical formulation is as follows:
$\mathbf{y}(\mathbf{p})=\sum^{K}_{k=1}\mathbf{w}_{k}\mathbf{m}_{k}\mathbf{x}(\mathbf{p}+\mathbf{p}_{k}+\Delta\mathbf{p}_{k}),$ (6)
where $\mathbf{p}$ denotes the current location in feature map, $\mathbf{w_{k}}$ and $\mathbf{p_{k}}$ denote the weight and initial offsets for the $\mathbf{k}$-th location, respectively. And $\mathbf{m_{k}}$ is a learnable modulation scalar, which helps to modulate the input feature amplitudes from different spatial locations. For example, $\mathbf{K}=9$ and $\mathbf{p_{k}}\in{(-1,-1),...,(1,1)}$ defines a $\mathbf{3\times 3}$ convolutional kernel with dilation 1.
The outputs from these two branches are concatenated and fed into a convolutional layer for feature fusion. This simple yet effective architecture successfully adapts the features extracted by the Mamba model, efficiently integrates multi-scale features, while endowing the entire decoder with deformable capabilities. Experiments on multiple datasets demonstrate both the effectiveness and efficiency of our proposed model.
4 Experiment Results
4.1 Datasets
To study the unification of wide-FoV segmentation, we conduct experiments on three 360° and two 180° datasets.
Stanford2D3D [1] consists of 1,413 indoor 360° images with a resolution of $512{\times}1024$, encompassing 13 distinct object categories.
Matterport3D [3] has 10,615 indoor 360° images. Following [53, 22], 7,829 of panoramas with a resolution of 1024×2048 are used for training and 772 for evaluation.
SynPASS [67] is a virtual outdoor 360° dataset from CARLA [13]. It contains 9,080 images with a resolution of $1024{\times}2048$, covering various weather conditions (cloudy, foggy, rainy, sunny) and illumination (daytime, nighttime).
WoodScape [64] consists of 10K 180° images captured by four surround view cameras. The semantic labels have 9 categories. Since the test split is not available, 8,234 released images are reallocated into training and evaluation sets with an 80% to 20% split.
SynWoodScape [48] is a synthetic version of the surround-view dataset. The released v0.1.1 version contains 2,000 180° surround-view fisheye RGB images. We split them into training and validation sets with an 80% to 20 % ratio.
4.2 Implementation Details
In this work, our models are trained and validated on three 360° datasets and two 180° datasets, covering both indoor and outdoor scenarios, as well as real and synthetic environments, using 4 A100 GPUs. For all experiments, we employed cross-entropy as the loss function. To validate the effectiveness of our framework, we deliberately avoided using any auxiliary losses. The models were optimized using AdamW with an initial learning rate of $6e^{-5}$ and weight decay of 0.01. For Stanford2D3D and SynWoodscape, we trained for 80K iterations with a batch size of 2 per GPU. For Matterport3D, SynPASS, and Woodscape, we extended the training to 160K iterations, maintaining the same batch size of 2 per GPU. The learning rate schedule consisted of a linear warm-up for the first 1.5K iterations, followed by a polynomial decay with a power of 0.9.
4.3 Quantitative Results
Given that previous works were largely constrained to specific Field-of-View (FoV) settings, which poses challenges for fair cross-dataset evaluation. Therefore, we selected representative state-of-the-art models as our baselines, including SegNeXt [21], which achieves superior performance on pinhole images, and 360SFUDA [73], which is specifically designed for panoramic scenarios. To ensure fair comparison, we carefully match the parameter counts across all implementations and conduct extensive evaluations on the aforementioned datasets with varying FoV characteristics.
Table 1: Results on 360° segmentation datasets of Stanford2D3D (S2D3D) fold-1 and 360FV-Matterport (MP3D). denotes our re-implementation.
  Method Backbone mIoU
S2D3D
MP3D
Tangent [14] (CVPR20) ResNet-101 45.6 -
HoHoNet [51] (CVPR21) ResNet-101 52.0 44.1
CBFC [74] (CVPR23) ResNet-101 52.2 -
Trans4PASS [66] (CVPR22) Trans4PASS 52.1 41.9
Trans4PASS+ [67] (PAMI24) Trans4PASS+ 54.0 42.6
SegFormer-B2 [58] (NeurIPS21) MiT-B2 51.9 45.5
360BEV [53] (WACV24) MiT-B2 54.3 46.3
SGAT4PASS [33] (IJCAI23) MiT-B2 56.4 -
360SFUDA [73] (CVPR24) MixT-B2 54.9 46.7
360SFUDA [73] (CVPR24) MixT-B3 56.9 47.9
SegNeXt [21] (NeurIPS22) MSCAN-B 57.0 48.4
VMamba [37] (NeurIPS24) VMamba-T 56.8 47.6
DMamba-M (Ours) VMamba-T 59.3 48.7
DMamba-T (Ours) VMamba-S 60.2 49.3
Results on 360° Stanford2D3D. Table 1 shows superior performance of our method compared to both panoramic [14, 51, 74, 66, 67, 53, 33, 73] and pinhole-based [21, 37] methods. Specifically, our model with the VMamba-T as backbone achieves a significant mIoU improvement of ${+}2.9\%$ to ${+}13.7\%$ over panoramic baselines.
Notably, compared to the VMamba-T + UperNet [57], our lightweighted DMamba-M with the same backbone achieves a ${+}2.5\%$ higher mIoU while reducing parameter count by 30% and FLOPs by 80%. Furthermore, our DMamba-T model, which has a comparable parameter count to VMamba-T + UperNet, demonstrates even better performance with a ${+}3.4\%$ mIoU improvement.
Table 2: Results on 360° segmentation dataset SynPASS. denotes our re-implementation.
  Method
mIoU
Building
Fense
Other
Pedestrian
Pole
RoadLine
Road
SideWalk
Vegetation
Vehicles
Wall
TrafficSign
Sky
Ground
Bridge
RailTrack
GroundRail
TrafficLight
Static
Dynamic
Water
Terrain
PVT-T [55] (CVM22) 32.37 74.83 19.94 00.24 21.82 13.15 62.59 93.14 49.09 67.27 46.44 01.69 09.63 96.09 00.18 02.64 08.81 61.11 14.09 12.04 00.99 05.05 51.33
PVT-S [55] (CVM22) 32.68 78.02 27.12 00.27 23.48 16.51 59.81 92.87 50.21 66.22 43.50 01.12 08.67 96.34 00.44 00.15 02.82 63.88 13.78 15.15 01.58 08.78 48.29
SegFormer-B1 [58] (NeurIPS21) 37.36 78.24 20.59 00.00 38.28 21.09 68.72 94.50 59.72 68.43 67.51 00.83 09.86 96.08 00.56 01.38 20.79 69.59 23.38 19.91 01.38 08.97 52.07
SegFormer-B2 [58] (NeurIPS21) 37.24 79.25 23.58 00.00 40.01 20.14 65.28 92.80 46.92 68.64 77.45 01.42 15.00 96.33 00.57 00.58 02.68 67.60 25.89 20.80 01.99 20.92 51.53
Trans4PASS-T [66] (CVPR22) 38.53 79.17 28.18 00.13 36.04 23.69 69.16 95.51 61.71 69.77 71.12 01.53 16.98 96.50 00.56 01.60 15.22 70.48 26.03 23.11 02.08 09.24 49.77
Trans4PASS-S [66] (CVPR22) 38.57 80.02 24.56 00.07 41.49 25.23 72.00 95.89 59.88 69.07 77.08 01.04 13.72 96.69 00.67 00.73 05.60 72.56 25.93 22.45 02.78 08.34 52.65
Trans4PASS+(T) [67] (PAMI24) 39.42 79.63 24.45 00.21 44.23 26.71 70.32 95.86 61.80 69.25 78.85 01.09 13.81 97.12 00.91 03.48 19.32 72.44 21.08 25.56 02.67 05.03 53.20
Trans4PASS+(S) [67] (PAMI24) 40.72 80.91 20.78 00.23 45.36 24.08 72.51 96.79 67.15 70.46 81.39 04.28 26.19 97.21 01.24 01.74 16.56 67.08 28.64 23.68 03.35 08.48 57.57
SegNeXt-B [21] (NeurIPS22) 39.90 82.62 16.06 00.12 43.21 26.15 71.69 97.10 60.64 67.56 82.56 00.32 24.43 97.23 02.03 00.00 43.23 47.90 31.61 26.50 03.63 10.56 42.68
360SFUDA-B2 [73] (CVPR24) 40.37 82.69 22.27 00.16 41.86 28.19 74.98 97.13 62.22 70.05 81.55 04.21 21.09 97.45 01.06 00.49 31.77 51.08 32.15 24.88 03.59 10.02 49.23
360SFUDA-B3 [73] (CVPR24) 41.33 82.91 32.33 00.23 41.87 27.21 75.13 97.37 62.90 69.34 82.95 01.66 26.03 97.56 00.95 00.10 41.39 64.02 32.30 26.74 03.48 16.76 48.14
DMamba-M(Ours) 41.72 83.19 21.52 00.21 48.15 32.20 75.8 97.51 61.00 67.74 81.86 01.06 31.30 94.47 01.37 00.68 29.63 49.41 36.35 30.99 04.08 20.54 45.27
DMamba-T(Ours) 42.78 84.24 22.57 00.25 49.20 33.25 76.85 98.56 62.05 68.79 82.91 02.11 32.35 98.52 02.42 01.73 30.68 50.46 37.40 32.04 05.85 21.59 46.32
Results on 360° Matterport3D.
While our DMamba-M demonstrates remarkable performance on Stanford2D3D dataset, the challenge becomes more pronounced on large-scale Matterport3D panoramic datasets, where we observe a relatively modest improvement margin of 0.3 mIoU over the previous SOTA model SegNeXt as shown in Table 1. Remarkably, our DMamba-T achieves more substantial gains of ${+}0.9\%$ and ${+}1.7\%$ mIoU compared to SegNeXt and VMamba-T respectively, with comparable model capacities.
Results on 360° SynPASS. Beyond real-world datasets, we evaluate the generalizability of our method on synthetic and outdoor scenes on the SynPASS dataset. It is worth noting that despite SegNeXt’s superior performance over 360SFUDA in real-world scenarios, it exhibits a substantial performance degradation on the synthetic dataset. This discrepancy underscores the sensitivity of certain architectures to distribution shifts. In contrast, our method, as shown in Table 2, achieves the state-of-the-art mIoU of 42.78%, surpassing all existing panoramic- and pinhole-based methods. Specifically, our DMamba-T overall achieves top scores on pedestrian, pole, traffic sign, static and dynamic. For traffic sign with slender properties, our model enjoys more than ${+}6.0\%$ mIoU.
Table 3: Results on 180° segmentation dataset WoodScape. denotes our re-implementation.
  Method Mutli-task Supervised mIoU
MeanTeacher [52] (NeurIPS17) 54.20
CPS [7] (CVPR21) 60.31
CPS with CutMix [7] (CVPR21) 62.47
FishSegSSL [43] (J. Imaging24) 64.81
OmniDet [31] (RA-L21) 72.50
    + DTP [20] (ECCV18) 75.80
    + GradNorm [8] (PMLR18) 75.90
    + Uncertainity [29] (CVPR18) 76.10
    + VarNorm [31] (RA-L21) 76.60
SegNeXt-B [21] (NeurIPS22) 82.32
360SFUDA-B2 [73] (CVPR24) 81.74
360SFUDA-B3 [73] (CVPR24) 82.43
DMamba-M (Ours) 83.21
DMamba-T (Ours) 83.90
Results on 180° WoodScape. In Table 3, we present a comparison of our method with other models using 180° fisheye inputs from the WoodScape dataset [64]. Semi-supervised methods such as MeanTeacher [52], CPS [7], and FishSegSSL [43] show lower performance due to the limited data. Compared to multi-task models, our proposed DMamba, designed with deformable settings, outperforms all variants of the OmniDet [31] model.
Although SegNeXt and 360SFUDA perform well on the WoodScape dataset, our method, with a more compact architecture still achieves a higher mIoU by ${+}0.9\%$ and ${+}0.8\%$, respectively. The DMamba-T variant further advances the state-of-the-art mIoU to 83.9%, confirming DMamba’s effectiveness in capturing heavily distorted object information in wide-FoV segmentation.
Table 4: Results on 180° segmentation dataset SynWoodScape. denotes our re-implementation.
             Method            mIoU
           SegNeXt-B [21] (NeurIPS22)            64.86
           360SFUDA-B2 [73] (CVPR24)            66.56
           360SFUDA-B3 [73] (CVPR24)            67.47
           DMamba-M (Ours)            69.20
           DMamba-T (Ours)            69.62
Results on 180° SynWoodScape. Due to the lack of baseline comparisons on SynWoodScape, we limited our evaluation to SegNeXt and 360SFUDA. Similar to the results on SynPASS, SegNeXt and 360SFUDA perform similarly on real 180° datasets, while the former exhibits a significant performance degradation on synthetic 180° data. This further confirms the challenge of achieving consistently high performance across diverse datasets. In comparison, our DMamba-M of similar scale outperforms them by ${+}4.36\%$ and ${+}1.73\%$ mIoU, respectively, achieving an improved segmentation accuracy of 69.20 mIoU on SynWoodScape.
These consistent improvements across all five wide-FoV datasets demonstrate the effectiveness of our proposed framework in unifying wide-FoV segmentation.
Figure 5: Visualization of the wide-FoV segmentation results. From left to right: one 360° and two 180° segmentation.
4.4 Qualitative Results
To demonstrate the improved performance of our model over Mamba-based models, we compared the visualization results of DMamba and VMamba on the Stanford2D3D and SynWoodScape. As discussed in Sec. 3.1 and  3.4, the scanning approach in Mamba-based models introduces biases in visual tasks that limit their effectiveness in dense prediction tasks. These biases increase randomness and reduce the model’s ability to capture local spatial relationships, which is particularly evident in the recognition and segmentation of smaller objects. As shown in Fig. 5, the baseline model struggles to accurately segment narrow and elongated areas such as door frames, vegetation contours, traffic light poles, and lane markings. In contrast, our proposed DMamba effectively addresses these challenges, significantly improving segmentation performance on small and elongated objects.
Table 5: Ablation study for analysis of backbones (BB) and decoders (Dec) on the 360° dataset Stanford2D3D.
[…]
BB
Dec
Total
BB
Dec
Total
mIoU
mAcc
ResNet-UperNet ResNet50 [23] UperHead [57] 44.3 23.5 40.5 64.0 22.9 250.7 273.6 53.34 62.42
ResNet-UperFormer ResNet50 [23] MusterHead [60] 14.3 23.5 203.1 226.6 22.9 211.5 234.4 54.39 63.27
ResNet-CGRSeg ResNet50 [23] CGRHead [41] 58.0 23.5 282.6 306.1 22.9 31.4 54.3 54.12 62.94
ResNet-DMamba (Ours) ResNet50 [23] DMambaHead 27.9 23.5 77.3 100.8 22.9 38.8 61.7 57.22 66.32
Swin-UperNet Swin-T [38] UperHead [57] 34.9 27.5 31.5 59.0 25.6 211.5 237.1 53.82 63.06
Swin-UperFormer Swin-T [38] MusterHead [60] 30.2 27.5 19.1 46.6 25.6 21.9 47.5 52.36 54.54
Swin-CGRSeg Swin-T [38] CGRHead [41] 52.2 27.5 40.6 68.1 25.6 5.5 31.1 54.26 62.90
Swin-DMamba (Ours) Swin-T [38] DMambaHead 37.3 27.5 11.2 38.7 25.6 6.5 32.1 55.09 65.30
VMamba-UperNet VMamba-T [37] UperHead [57] 35.0 29.5 31.5 61.0 25.3 206.9 232.2 56.80 65.01
VMamba-UperFormer VMamba-T [37] MusterHead [60] 29.8 29.5 19.1 48.6 25.3 21.4 46.7 58.12 66.97
VMamba-CGRSeg VMamba-T [37] CGRHead [41] 42.2 29.5 40.5 70.0 25.3 5.0 30.3 56.45 65.07
DMamba-M (Ours) VMamba-T [37] DMambaHead 40.0 29.5 11.1 40.6 25.3 6.0 31.3 59.30 68.99
DMamba-T (Ours) VMamba-S [37] DMambaHead 17.9 49.4 11.1 60.5 44.7 6.0 50.7 60.18 71.68
4.5 Ablation Study
Effect of Backbone. In Table 5, we ablate three different kinds of commonly-used model backbones, namely CNN-based ResNet [23], ViT-based Swin Transformer [38] and Mamba-based VMamba [37]. For the sake of a fair comparison, we choose the backbones with a comparable number of learnable parameters and FLOPs. According to Table 5 it can be observed that the models equipped with VMamba [37] outperform other models consistently while CNN-based ResNet50 and ViT-based Swin Transformer achieve similar performance. This conclusion can be validated by employing the same decoder, such as UperHead [57], MusterHead [60], CGRHead [41], and our proposed DMambaHead, along with a consistent model backbone to perform the segmentation task. Notably, the combination of the VMamba [37] backbone with DMambaHead, resulting in our Deformable Mamba model, demonstrates superior performance compared to other variant combinations by a maximum of over $+6.9\%$ mIoU.
Effect of Decoder. Apart from the backbones, we also ablate the decoders. For every backbone, we conduct four experiments with different decoders, i.e., UperHead [57], MusterHead [60], CGRHead [41], and our proposed DMambaHead. All results are reported in Table 5. Using the same model backbone, our proposed DMambaHead outperforms other widely-used decoders by a $+3\%$ mIoU while a complexity degradation of $+20\%$, making the Deformable Mamba and its variants extremely practical in real-time wide-FoV accurate segmentation. It’s worth noting that the DMambaHead with VMamba-T [37] backbone achieves a drop of FLOPs from $232.2$G to $31.3$G, a drop of the learnable parameters from $61$M to $40.6$M while the performance of Deformable Mamba increases from a $+56.8\%$ to $+59.3\%$ mIoU compared with the VMamba-T [37] and UperHead [57] combination. The reason behind this is that our proposed DMambaHead is specifically designed for wide-FoV segmentation. Our deformable design enables the model to focus on the distortion introduced by the wide-FoV fisheye or panoramic images. This ablation proves that not only the VMamba [37] backbone but also the proposed DMambaHead contributes to the superior performance and efficiency of the Deformable Mamba model.
(a) Results on Matterport3D.
(b) Results on Stanford2D3D.
Figure 6: Analysis of segmentation with different wide FoVs.
Analysis of FoVs. To further analyze the wide-FoV segmentation, we conduct omni-directional analysis. As shown in Fig. 6, there are six different FoVs, ranging from 60° to 360°. Compared to the baseline VMamba [37], our DMamba methods obtain relative better results on all different FoVs. Note that the 60° performance of all methods are lower than other degrees on the Stanford2D3D dataset. One reason is that the object distribution and room layout in the 60° direction are different between Stanford2D3D and the Matterport3D datasets. Nonetheless, our Mini- and Tiny-version methods show consistent improvements over the baseline on both datasets in different FoVs.
5 Conclusion
In this work, we are the first to explore the unification of the wide field of view (FoV) segmentation tasks, including 180° and 360° image segmentation. Most existing models are tailored for pinhole camera images, making it challenging to extend their success to wide-FoV domains due to the severe image distortion and object deformation. Unlike previous approaches that require a specific architecture for each FoV or dataset, we propose a unified framework, Deformable Mamba, built on the emerging State Space Model (SSM) mechanism. Our framework is constructed with a flexible deformable decoder, which can be easily adapted to various models and datasets. Consistent improvements across five diverse wide-FoV datasets, ranging from indoor to outdoor, from synthetic to real-world scenes, demonstrate that our unified framework effectively addresses the unique challenges of wide-FoV segmentation.
Limitations and Future Work. Using large language models (LLMs) for unifying the wide-FoV segmentation is a potential solution but not discussed in this work, while LLMs can further include the segmentation methods that use vision-language architectures. Besides, cross-domain evaluation, e.g., transferring from a synthetic to a real wide-FoV scene, or from a narrow-FoV camera to a wide-FoV camera, will be further explored in our future work.